OpenStudy (anonymous):

Find the area of region consisting of points (x,y) which satisfy sin(x+y)>=0 and x^2+y^2<=100

OpenStudy (anonymous):

x+y=npi y=npi-x

OpenStudy (anonymous):

oh pellet..sorry

OpenStudy (anonymous):

now how will u find the area of this??

OpenStudy (anonymous):

explain more please.. :)

OpenStudy (anonymous):

what?? "Since sin(x+y)>=0 is only true for quadrants 1 and 2. You only have the top half of the circle" what about x=5,y=-5? its in 4 quad

OpenStudy (amistre64):

Find the area of region consisting of points (x,y) which satisfy: sin(x+y) >= 0 x^2+y^2 <= 100 the sin function itself is between -1 and 1 and x^2+y^2 = 10 is a circle of radius 10

OpenStudy (amistre64):

lol.... thats a cute picture :)

OpenStudy (anonymous):

hmmm

OpenStudy (amistre64):

could we possibly convert the x^2+y^2 into a polar equation? so that we can compare the 2?

OpenStudy (amistre64):

sin(x+y) >= 0 sin(x)cos(y) + sin(y)cos(x) >= 0 might help?

OpenStudy (nikvist):

Sum of all regions is semicircle

OpenStudy (nikvist):

\[S=\frac{1}{2}S_{circle}=\frac{1}{2}10^2\pi=50\pi\]

OpenStudy (anonymous):

ya...but its a semicircle how will u prove?